# Mapping walnut water stress with high resolution multispectral UAV imagery and machine learning

### 30 Dec 2023 v1

@@ Abstract
Effective monitoring of walnut water status and stress level across the whole orchard is an essential step towards precision irrigation management of walnuts, a significant crop in California. This study presents a machine learning approach using Random Forest (RF) models to map stem water potential (SWP) by integrating high-resolution multispectral remote sensing imagery from Unmanned Aerial Vehicle (UAV) flights with weather data. From 2017 to 2018, five flights of an UAV equipped with a seven-band multispectral camera were conducted over a commercial walnut orchard, paired with concurrent ground measurements of sampled walnut plants. The RF regression model, utilizing vegetation indices derived from orthomosaiced UAV imagery and weather data, effectively estimated ground-measured SWPs, achieving an $R^{2}$ of 0.63 and a mean absolute error (MAE) of 0.80 bars. The integration of weather data was particularly crucial for consolidating data across various flight dates. Significant variables for SWP estimation included wind speed and vegetation indices such as NDVI, NDRE, and PSRI. A reduced RF model excluding red-edge indices of NDRE and PSRI, demonstrated slightly reduced accuracy ($R^{2}$ = 0.54). Additionally, the RF classification model predicted water stress levels in walnut trees with 85% accuracy, surpassing the 80% accuracy of the reduced classification model. The results affirm the efficacy of UAV-based multispectral imaging combined with machine learning, incorporating thermal data, NDVI, red-edge indices, and weather data, in walnut water stress estimation and assessment. This methodology offers a scalable, cost-effective tool for data-driven precision irrigation management at an individual plant level in walnut orchards.
@@
pagerange: Mapping walnut water stress with high resolution multispectral UAV imagery and machine learningMapping walnut water stress with high resolution multispectral UAV imagery and machine learning
@@ 1 Introduction
California’s Central Valley is a pivotal region for walnut production, contributing over 99% of the walnuts grown in the U.S. and supplying about half of the global walnut trade. This makes walnuts one of the most economically significant specialty crops in California, with an annual retail value exceeding \$15.4 billion (Federal Register 2023). However, the region’s recent severe drought conditions have compelled growers to rely heavily on groundwater pumping, leading to alarmingly rapid rates of groundwater depletion (Famiglietti, 2014). In response to these challenges, the optimization of water management practices has become a critical necessity for crop producers, especially given the escalating water demands tied to the expansion of specialty crops in the Central Valley. Timely and accurate mapping of water status across entire blocks is essential for implementing adaptive and precision irrigation management at both zonal and individual walnut plant levels, thereby enhancing water use efficiency. This is particularly crucial when initiating irrigation (Intrigliolo & Castel, 2007). Conventionally, walnut plant physiological water status is gauged through stem and/or leaf water potentials using pressure chambers (Choné et al., 2001; Williams & Araujo, 2002). However, this approach, being labor-intensive and typically limited to a small number of sample walnut plants, does not effectively support the mapping of within-field variability or the monitoring of temporal changes in orchard blocks.
@@
Meanwhile, various machine learning techniques have been applied to predict water stress dynamics in crops (Romero Ponce et al., 2018), with some advanced studies incorporating weather data to improve accuracy (Tang et al., 2023, 2022). Deep learning methods, including convolutional neural networks (CNN) and Inception-Resnet, have been used to classify multispectral imaging (MSI) of agricultural fields. These methods leverage pixelwise spectral data, spectral-spatial features, and sometimes additional principal components (Hsieh & Kiang, 2020; Liyuan et al., 2019). Typically, deep learning models require thousands of training samples, and various data augmentation techniques for these models are discussed in Illarionova et al. (2021).
@@ 2.5 Selection of Vegetation Indices
Physiological research has established that weather conditions, such as air temperature, relative humidity, and vapor pressure deficit (VPD), play a crucial role in regulating plant stomatal closure and atmospheric water exchange, thereby directly influencing plant water status (Garnier & Berger, 1985). In light of this, models that integrate multiple VIs from UAV imagery with weather variables have shown promising results in monitoring plant water stress. An example by Tang et al. (2022) successfully incorporated weather data into such a model, achieving accuracy in plant water stress prediction.
@@ 3.1 Full RF Regression Model Based on Single Flight Date Data
The performance of the RF regression model was first evaluated using data collected on individual flight dates. Given that weather conditions remain constant within a single flight date, the model inputs were reduced to thermal data and vegetation indices NDVI, NDRE, and PSRI. Figure 7 presents the results of this analysis. Mean $R^{2}$ scores for the different flight dates were -0.28 (+/- 0.35) for 07/24/2017, -1.89 (+/- 4.03) for 08/22/2017, -0.17 (+/- 0.79) for 07/09/2018, and -2.01 (+/- 3.40) for 07/27/2018, respectively. These values indicate the model’s varying levels of predictive accuracy on different dates. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 8, 9, 10 and 11.
@@ 3.2 Full RF Regression Model Based on Data from All Flight Dates
As depicted in Figure 12a, this comprehensive approach resulted in a mean $R^{2}$ value of 0.63 (+/- 0.34), (MAE: 0.80, MSE: 1.03), a significant increase compared to the single-date model results, highlighting a substantial enhancement in predictive accuracy. Furthermore, the right plot of the figure elucidates the variable importance ranking. It identifies the most impactful factors in descending order of influence: wind, VPD, thermal data, NDVI, NDRE, PSRI, and air temperature. This ranking offers valuable insights into the relative contribution of each variable to the performance of the model. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 13.
@@ 3.3 NoRedEdge RF Regression Model Based on Single Flight Date Data
The study progressed to assess a simplified RF model, specifically excluding the red-edge indices (NDRE and PSRI). This streamlined model, tailored for data from each individual flight date, incorporated only thermal data and NDVI as its input variables. The analysis results are presented in Figure 14. Mean $R^{2}$ scores for the different flight dates were -0.45 (+/- 0.60) for 07/24/2017, -2.77 (+/- 5.33) for 08/22/2017, -0.19 (+/- 0.55) for 07/09/2018, and -1.85 (+/- 2.96) for 07/27/2018, respectively. Comparatively lower than the $R^{2}$ values of the full RF model, these results demonstrate that excluding red-edge indices will reduce the model’s predictive accuracy. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 15, 16, 17 and 18.
@@ 3.4 NoRedEdge RF Regression Model Based on Data from All Flight Dates
The results, as shown in Figure 19a, reveal a mean $R^{2}$ score of 0.54 (+/- 0.38). This represents a notable improvement over the single-date models, yet it is marginally lower than the $R^{2}$ value of the full RF model. The right plot of the figure illustrates the variable importance, ranking the factors in descending order of influence: wind, thermal data, NDVI, VPD, and air temperature. This ranking provides insights into the contributions of these variables to the model’s predictive capability. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 20.

### 10 Jan 2024 v2

# vegetation index, unmanned aerial vehicles, multispectral imagery, stem water potential, water stress, machine learning, random forest

@@ Abstract
Effective monitoring of walnut water status and stress level across the whole orchard is an essential step towards precision irrigation management of walnuts, a significant crop in California. This study presents a machine learning approach using Random Forest (RF) models to map stem water potential (SWP) by integrating high-resolution multispectral remote sensing imagery from Unmanned Aerial Vehicle (UAV) flights with weather data. From 2017 to 2018, five flights of an UAV equipped with a seven-band multispectral camera were conducted over a commercial walnut orchard, paired with concurrent ground measurements of sampled walnut plants. The RF regression model, utilizing vegetation indices derived from orthomosaiced UAV imagery and weather data, effectively estimated ground-measured SWPs, achieving an $R^{2}$ of 0.70 and a mean absolute error (MAE) of 0.80 bars. The integration of weather data was particularly crucial for consolidating data across various flight dates. Significant variables for SWP estimation included wind speed and vegetation indices such as NDVI, NDRE, and PSRI. As a comparison, a reduced RF model excluding red-edge indices of NDRE and PSRI, demonstrated slightly reduced accuracy ($R^{2}$ = 0.63). Subsequently, a random-forest classification model was developed to categorize the severity of walnut plant water stress to three distinct stress levels: low stress, moderate stress, and severe stress.The RF classification model predicted water stress levels in walnut trees with 85% accuracy, surpassing the 80% accuracy of the reduced classification model. The results affirm the efficacy of UAV-based multispectral imaging combined with machine learning, incorporating thermal data, NDVI, red-edge indices, and weather data (wind, temperature, VPD), in walnut water stress estimation and assessment. This methodology enables a scalable, cost-effective tool for data-driven precision irrigation management at an individual plant level in walnut orchards. It highlights the promising capability to replace manual, labor-intensive measurements with precise, targeted irrigation, marking a significant advancement in refining and optimizing irrigation strategies.
@@
[…]
@@ 1 Introduction
California’s Central Valley is a pivotal region for walnut production, contributing over 99% of the walnuts grown in the U.S. and supplying about half of the global walnut trade. This makes walnuts one of the most economically significant specialty crops in California, with an annual retail value exceeding \$15.4 billion (Federal Register 2023). However, the region’s recent severe drought conditions have compelled growers to rely heavily on groundwater pumping, leading to alarmingly rapid rates of groundwater depletion (Famiglietti, 2014). In response to these challenges, the optimization of water management practices has become a critical necessity for crop producers, especially given the escalating water demands tied to the expansion of specialty crops in the Central Valley. Timely and accurate mapping of water status across entire blocks is essential for implementing adaptive and precision irrigation management at both zonal and individual walnut plant levels, thereby enhancing water use efficiency. This is particularly crucial when initiating irrigation (Intrigliolo and Castel, 2007). Conventionally, walnut plant physiological water status is gauged through stem and/or leaf water potentials using pressure chambers (Choné et al., 2001; Williams and Araujo, 2002). However, this approach, being labor-intensive and typically limited to a small number of sample walnut plants, does not effectively support the mapping of within-field variability or the monitoring of temporal changes in orchard blocks.
@@
Meanwhile, various machine learning techniques have been applied to predict water stress dynamics in crops (Romero Ponce et al., 2018), with some advanced studies incorporating weather data to improve accuracy (Tang et al., 2023, 2022). Deep learning methods, including convolutional neural networks (CNN) and Inception-Resnet, have been used to classify multispectral imaging (MSI) of agricultural fields. These methods leverage pixelwise spectral data, spectral-spatial features, and sometimes additional principal components (Hsieh and Kiang, 2020; Liyuan et al., 2019). Typically, deep learning models require thousands of training samples, and various data augmentation techniques for these models are discussed in Illarionova et al. (2021).
@@ 2.5 Selection of Vegetation Indices
Physiological research has established that weather conditions, such as air temperature, relative humidity, and vapor pressure deficit (VPD), play a crucial role in regulating plant stomatal closure and atmospheric water exchange, thereby directly influencing plant water status (Garnier and Berger, 1985). In light of this, models that integrate multiple VIs from UAV imagery with weather variables have shown promising results in monitoring plant water stress. An example by Tang et al. (2022) successfully incorporated weather data into such a model, achieving accuracy in plant water stress prediction.
@@ 3.1 Full RF Regression Model Based on Single Flight Date Data
The performance of the RF regression model was first evaluated using data collected on individual flight dates. Given that weather conditions remain constant within a single flight date, the model inputs were reduced to thermal data and vegetation indices NDVI, NDRE, and PSRI. Figure 7 presents the results of this analysis. Mean $R^{2}$ scores for the different flight dates were 0.08 (+/- 0.51) for 07/24/2017, -1.89 (+/- 3.40) for 08/22/2017, 0.17 (+/- 0.33) for 07/09/2018, and -0.03 (+/- 0.52) for 07/27/2018, respectively. These values indicate the model’s varying levels of predictive accuracy on different dates. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 9, 9, 11 and 11.
@@ 3.2 Full RF Regression Model Based on Data from All Flight Dates
As depicted in Figure 12, this comprehensive approach resulted in a mean $R^{2}$ value of 0.65 (+/- 0.15), a significant increase compared to the single-date model results, highlighting a substantial enhancement in predictive accuracy. Furthermore, the right plot of the figure elucidates the variable importance ranking. It identifies the most impactful factors in descending order of influence: wind, VPD, thermal data, NDVI, NDRE, PSRI, and air temperature. This ranking offers valuable insights into the relative contribution of each variable to the performance of the model. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 13.
@@ 3.3 NoRedEdge RF Regression Model Based on Single Flight Date Data
The study progressed to assess a simplified RF model, specifically excluding the red-edge indices (NDRE and PSRI). This streamlined model, tailored for data from each individual flight date, incorporated only thermal data and NDVI as its input variables. The analysis results are presented in Figure 14. Mean $R^{2}$ scores for the different flight dates were -0.42 (+/- 0.87) for 07/24/2017, -3.50 (+/- 6.72) for 08/22/2017, -0.76 (+/- 1.42) for 07/09/2018, and -1.05 (+/- 1.06) for 07/27/2018, respectively. Comparatively lower than the $R^{2}$ values of the full RF model, these results demonstrate that excluding red-edge indices will reduce the model’s predictive accuracy. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 16, 16, 18 and 18.
@@ 3.4 NoRedEdge RF Regression Model Based on Data from All Flight Dates
The results, as shown in Figure 19, reveal a mean $R^{2}$ score of 0.59 (+/- 0.17). This represents a notable improvement over the single-date models, yet it is marginally lower than the $R^{2}$ value of the full RF model. The right plot of the figure illustrates the variable importance, ranking the factors in descending order of influence: wind, thermal data, NDVI, VPD, and air temperature. This ranking provides insights into the contributions of these variables to the model’s predictive capability. Furthermore, to illustrate the influence of each variable on SWP, partial dependence plots (PDPs) are presented in Figure 20.